The width of a rectangular lot is 75% of its length. If the perimeter is 1050 meters, then what are the length and the width?
Answer by tutorcecilia(2144)

The width of a rectangular lot is 75% of its length. If the perimeter is 1050 meters, then what are the length and the width?
P = 2(length + width) [Formula for the perimeter]
The width = (.75)(length)
.
1050 = 2(length + (.75)(length)) [Plug the values into the formula]
1050 = 2(x + .75x)
1050/2 = 2/2(x + .75x) [Solve for x]
525 = x + .75x
525 = 1.75x
525/1.75 = 1.75/1.75x
300=x = the length
.
(75% of 300) [The width is 75% of the length]
(.75)(300) = 225 [The width]
.
Check by plugging (x=300) back into the equation
.
P = 2(length + width) [Formula for the perimeter]
1050 = 2(300 + (.75)(300))
1050 = 2( 300 + 225)
1050 = 2(525)
1050 = 1050 [Checks out]
So, the length is 300 and the width is 225 or (75% of 300 = 225)